# Riemannian Variational Flow Matching for Material and Protein Design

### v2

@@ Abstract
We present Riemannian Gaussian Variational Flow Matching (RG-VFM), a geometric extension of Variational Flow Matching (VFM) for generative modeling on manifolds. In Euclidean space, predicting endpoints (VFM), velocities (FM), or noise (diffusion) are largely equivalent due to affine interpolations. On curved manifolds this equivalence breaks down, and we hypothesize that endpoint prediction provides a stronger learning signal by directly minimizing geodesic distances. Building on this insight, we derive a variational flow matching objective based on Riemannian Gaussian distributions, applicable to manifolds with closed-form geodesics. We formally analyze its relationship to Riemannian Flow Matching (RFM), exposing that the RFM objective lacks a curvature-dependent penalty – encoded via Jacobi fields – that is naturally present in RG-VFM. Experiments on synthetic spherical and hyperbolic benchmarks, as well as real-world tasks in material and protein generation, demonstrate that RG-VFM more effectively captures manifold structure and improves downstream performance over Euclidean and velocity-based baselines.
@@ 1 Introduction
Recent developments have extended flow matching in two key directions. Variational Flow Matching (VFM) (Eijkelboom et al., 2024) reframes the problem as posterior inference over trajectories, providing a probabilistic perspective with flexible modeling choices. In parallel, Riemannian Flow Matching (RFM) (Chen & Lipman, 2024) has shown how incorporating non-Euclidean geometry can improve modeling of distributions supported on manifolds.
At a high level, generative models can be trained by predicting an endpoint (as in VFM), a velocity (as in FM/RFM), or noise (as in diffusion). In Euclidean space, these choices are largely equivalent up to affine transformations, since predicting noise, score, velocity field or endpoints are interchangeable parameterizations of the same training signal (Vuong et al., 2025; Lipman et al., 2023; Eijkelboom et al., 2024). On curved manifolds, however, this equivalence breaks down: tangent spaces differ across points and curvature introduces higher-order deviations. We hypothesize that predicting endpoints provides a stronger learning signal in the geometric setting, as it directly minimizes geodesic distances and aligns training with intrinsic manifold structure.
@@
This geometric extension is particularly relevant for biological and chemical domains where intrinsic geometric structure governs the data. Recently, generative models have been extensively applied to material discovery (Jiao et al., 2023; Fu et al., 2023; Kim et al., 2024) and to the generation of large biomolecules such as protein backbones (Guo et al., 2025; Yue et al., 2025; Yim et al., 2023a; b). These applications highlight that data often live on heterogeneous manifolds: Euclidean space for atomic coordinates, rotation groups $\mathrm{SO}(3)$ for orientations, and other structured domains. Early works in crystal generation, such as Jiao et al. (2023), focus purely on Euclidean parameters without explicitly modeling rotational degrees of freedom. In contrast, recent methods for metal-organic frameworks (MOFs) and proteins (Yim et al., 2023a; Yue et al., 2025; Kim et al., 2024; Guo et al., 2025) adopt a mixed approach where Euclidean parameters (e.g., positions) are modeled with standard FM while non-Euclidean parameters (e.g., rotations) are modeled with RFM. These methods lack a fully variational treatment across both parameter types, and we address this gap by demonstrating the benefits of our geometric variational approach on these applications.
@@
In this paper, we develop Riemannian Gaussian Variational Flow Matching (RG-VFM), which bridges the variational and geometric extensions of flow matching by extending VFM to Riemannian manifolds with closed-form metrics. Our contributions are threefold:
@@ 2 Background
Riemannian Flow Matching (RFM) (Chen & Lipman, 2024) extends FM to Riemannian manifolds. Given a smooth Riemannian manifold $\mathcal{M}$ with closed-form geodesics and metric $\mathbf{g}$, RFM learns a vector field $v_{t}$:
@@
Variational Flow Matching (VFM) (Eijkelboom et al., 2024) reformulates FM by introducing a variational distribution $q_{t}^{\theta}(x_{1}\mid x)$ to approximate the unknown posterior $p_{t}(x_{1}\mid x)$, where the learned velocity $v_{t}^{\theta}$ is expressed as the expectation of the condition velocity under this variational approximation over trajectories. Then, the VFM objective is to minimize the KL divergence between joint distributions, i.e.:
@@
$\mathcal{L}_{\mathrm{VFM}}(\theta)=-\mathbb{E}_{t,x_{1},x}\left[\sum_{d=1}^{D}\log q_{t}^{\theta}(x_{1}^{d}\mid x)\right],\text{ e.g. }\mathcal{L}_{\mathrm{VFM}}(\theta)=\mathbb{E}_{t,x_{1},x}\left[\,\|\mu_{t}^{\theta}(x)-x_{1}\|^{2}\right],$ (5)
if $q_{t}^{\theta}$ is Gaussian, relating VFM directly back to FM (see Eijkelboom et al. (2024) for details). A key feature of VFM is its flexibility in choosing $q_{t}^{\theta}$, as different choices allow adaptation to various geometries and data types, improving efficiency and expressiveness. VFM has demonstrated advantages over standard FM in discrete domains (e.g., CatFlow) and has been extended to mixed data modalities (Guzmán-Cordero et al., 2025) as well as molecular generation tasks (Eijkelboom et al., 2025; Sakalyan et al., ). A key strength of the variational formulation is its flexibility: problem-specific constraints can be incorporated directly into the objective. For instance, censored flow matching for sea-ice forecasting enforces physical bounds such as non-negative ice thickness through the variational loss (Finn et al., 2025).
@@ 3 Riemannian Gaussian Variational Flow Matching
  • Flexibility on the support of the distribution: the prior $p_{0}$ can be defined either on $\mathcal{M}$ (intrinsic) or in the ambient Euclidean space (extrinsic), while vanilla RFM only supports the intrinsic viewpoint. The extrinsic framework maintains the simplicity and efficiency of a linear flow, avoiding the need for the manifold’s exponential and logarithmic maps, while encoding more geometric information than purely Euclidean methods (section 3.1).
@@ 3.1 The Riemannian Gaussian VFM Objective
To extend VFM to the geometric case, one first needs to define a relevant variational posterior with support over the manifold. Let $\mathcal{M}$ be a Riemannian manifold with metric $\mathbf{g}$. The Riemannian Gaussian (RG) distribution (Pennec, 2006) is defined as the maximum–entropy distribution specified by its mean value and covariance, formally
@@
In the Euclidean Gaussian VFM case, this setting reduces to a straightforward mean squared error optimization, so it is natural to wonder whether a similar simplification holds here. In fact, such a simplification exists under two assumptions: (1) the manifold is homogeneous – that is, any point can be transformed into any other by a distance-preserving symmetry (a formal definition is provided in section C.1); and (2) we have access to a closed-form expression for its geodesics. Formally, the following holds (see section D.1 for details):
@@
Minimizing this loss is equivalent to computing the Fréchet mean of the distribution, that is: $\mu^{\star}=\operatorname*{arg\,min}_{\mu_{\theta}\in\mathcal{M}}\mathbb{E}_{x_{1}}\left[\text{dist}_{\mathbf{g}}(x_{1},\mu_{\theta})^{2}\right]$, averaged over the generative steps $t$ and samples $x$. In other words, the point $\mu_{\theta}$ minimizes the expected squared geodesic distance to the target (Fréchet, 1948). Intuitively, this can be viewed as a generalization of the mean squared error from the Euclidean setting to a Riemannian framework.
@@
The extrinsic variant RG-VFM-$\mathbb{R}^{n}$ thus learns a simple linear flow while retaining a geometry-aware loss, whereas the intrinsic variant RG-VFM-$\mathcal{M}$ mirrors the RFM setup but differs in its loss definition. Indeed, Vanilla RFM also requires $\operatorname{supp}(p_{0})\subseteq\mathcal{M}$ because its vector fields depend on the manifold’s intrinsic geometry. Because of these different frameworks, direct comparison is only meaningful between RG-VFM-$\mathcal{M}$ and RFM, a comparison we present in the next section.
@@ 4.1 Jacobi Field formulation of the Flow Matching Objectives
Borrowing the notations from (Chen & Lipman, 2024), we denote the target velocity $v^{0}=u_{t}(x\mid x_{1})$, the predicted velocity $v^{1}=v_{t}^{\theta}(x)$, and their respective endpoints $\gamma_{0}(1)=x_{1}$ and $\gamma_{1}(1)=\mu_{t}^{\theta}(x)$. The losses can be formulated in the Jacobi field framework with the following:
@@
Consider a Jacobi field $J(\tau):=\partial_{s}\alpha(s,\tau)\big|_{s=0}$ as defined in definition 4.1. We denote ${\mathcal{L}}_{\mathrm{RFM}}$ the loss function of the (vanilla) Riemannian Flow Matching (Chen & Lipman, 2024), and ${\mathcal{L}}_{\mathrm{RG\text{-}VFM}}$ the loss function for our proposed Riemannian Variational Flow Matching. Then the following equalities hold:
@@ 4.2 Relation between RG-VFM and RFM Objectives
The curvature functional $\mathcal{C}$ captures how the manifold’s geometry affects the loss comparison, encoding the first- and second-order effects of curvature on geodesic deviation. Thus, RG-VFM captures the full geometric structure through the exact Jacobi field $J(1)$, while RFM uses only the linear approximation $D_{\tau}J(0)$. This lack of curvature information results in weaker, less precise supervision in directing the flow toward the actual endpoint, leading in practice to RG-VFM learning the signal more effectively than RFM. Special cases are:
@@
  • In spaces of constant curvature (e.g. hyperspheres or hyperbolic spaces) $\nabla R=0$. In this setting, we can restate the result of proposition 4.3 in terms of the constant sectional curvature $K$. The formulation and proof are given in corollary D.1.
In summary, we introduced RG-VFM as an alternative to RFM for learning a velocity field on a manifold, providing a variational formulation whose objective fully captures higher-order curvature effects, unlike RFM. This results in generally different objectives on curved manifolds. In Euclidean space, however, the RFM objective reduces to CFM, while RG-VFM reduces to VFM (assuming a Euclidean Gaussian posterior rather than Riemannian), and the CFM and VFM objectives become equivalent under appropriate normalization. These relations are schematized in fig. 1 and fig. 2(b).
@@ 5.1 Curvature Effects in Synthetic Data
We conduct two sets of experiments: we (1) compare the extrinsic models in their ability to capture the correct geometry – assessed by the distance of the generated samples to the data manifold – and (2) evaluate vanilla versus variational models in reproducing the target distribution. For this last point, the evaluation is based on metrics such as Coverage (% of generated points falling within the desired checkerboard area) and Classifier 2-Sample Tests (C2ST) metric (Lopez-Paz & Oquab, 2016; Dalmasso et al., 2020; Lueckmann et al., 2021). The C2ST technique utilizes a neural network classifier to separate true samples from generated ones, where a score of 0.5 indicates the distributions are indistinguishable to the classifier, while scores approaching 1 suggest the distributions are easily separable. Additional experimental details are provided in appendix E.

### v3

@@ Abstract
We present Riemannian Gaussian Variational Flow Matching (RG-VFM), a geometric extension of Variational Flow Matching (VFM) for generative modeling on manifolds. Motivated by the benefits of VFM, we derive a variational flow matching objective for manifolds with closed-form geodesics based on Riemannian Gaussian distributions. Crucially, in Euclidean space, predicting endpoints (VFM), velocities (FM), or noise (diffusion) is largely equivalent due to affine interpolations. However, on curved manifolds this equivalence breaks down. We formally analyze the relationship between our model and Riemannian Flow Matching (RFM), revealing that the RFM objective lacks a curvature-dependent penalty – encoded via Jacobi fields – that is naturally present in RG-VFM. Based on this relationship, we hypothesize that endpoint prediction provides a stronger learning signal by directly minimizing geodesic distances. Experiments on synthetic spherical and hyperbolic benchmarks, as well as real-world tasks in material and protein generation, demonstrate that RG-VFM more effectively captures manifold structure and improves downstream performance over Euclidean and velocity-based baselines.
@@ 1 Introduction
Recent developments have extended flow matching in two key directions. Variational Flow Matching (VFM) (Eijkelboom et al., 2024) reframes the problem as posterior inference over trajectories, providing a probabilistic perspective with flexible modeling choices. In parallel, Riemannian Flow Matching (RFM) (Chen and Lipman, 2024) has shown how incorporating non-Euclidean geometry can improve modeling of distributions supported on manifolds.
VFM has demonstrated advantages over standard FM in discrete domains (e.g., CatFlow) and has been extended to mixed data modalities (Guzmán-Cordero et al., 2025) as well as molecular generation tasks (Eijkelboom et al., 2025; Sakalyan et al., ). A key strength of the variational formulation is its flexibility: problem-specific constraints can be incorporated directly into the objective. For instance, censored flow matching for sea-ice forecasting enforces physical bounds such as non-negative ice thickness through the variational loss (Finn et al., 2025). Similarly, when the target distributions live on a Riemannian manifold, the objective can be formulated to respect the geometry of the data support, which is the setting we study in this paper.
@@
The geometric extension is particularly relevant for biological and chemical domains where intrinsic geometric structure governs the data. Recently, generative models have been extensively applied to material discovery (Jiao et al., 2023; Fu et al., 2023; Kim et al., 2024) and to the generation of large biomolecules such as protein backbones (Guo et al., 2025; Yue et al., 2025; Yim et al., 2023a; b). These applications highlight that data often live on heterogeneous manifolds: Euclidean space for atomic coordinates, rotation groups $\mathrm{SO}(3)$ for orientations, and other structured domains. Early works in crystal generation, such as Jiao et al. (2023), focus purely on Euclidean parameters without explicitly modeling rotational degrees of freedom. In contrast, recent methods for metal-organic frameworks (MOFs) and proteins (Yim et al., 2023a; Yue et al., 2025; Kim et al., 2024; Guo et al., 2025) adopt a mixed approach where Euclidean parameters (e.g., positions) are modeled with standard FM while non-Euclidean parameters (e.g., rotations) are modeled with RFM. These methods lack a fully variational treatment across both parameter types, and we address this gap by demonstrating the benefits of our geometric variational approach on these applications.
When extending from Euclidean space to general Riemannian manifolds, it becomes unclear how different generative objectives relate. In Euclidean space, training a generative model by predicting an endpoint (VFM), a velocity (FM), or noise (diffusion) is largely equivalent up to affine transformations, since these quantities – noise, score, velocity field, and endpoints – are interchangeable parameterizations of the same training signal (Vuong et al., 2025; Lipman et al., 2023; Eijkelboom et al., 2024). On curved manifolds, however, this near-equivalence breaks down: tangent spaces vary across points and curvature introduces higher-order deviations, preventing any explicit closed-form relation between the velocity-based and endpoint-based perspectives. This raises the questions: if the endpoint and velocity-based perspectives are no longer equivalent on curved manifolds, how do they differ, and which one is preferable?
@@
In this paper, we develop Riemannian Gaussian Variational Flow Matching (RG-VFM), which extends VFM to Riemannian manifolds with closed-form metrics, thereby bridging the variational and geometric extensions of flow matching. Our contributions are threefold:
@@ 2 Background
Riemannian Flow Matching (RFM) (Chen and Lipman, 2024) extends FM to Riemannian manifolds. Given a smooth Riemannian manifold $\mathcal{M}$ with closed-form geodesics and metric $\mathbf{g}$, RFM learns a vector field $v_{t}$:
@@
Variational Flow Matching (VFM) (Eijkelboom et al., 2024) reformulates FM by introducing a variational distribution $q_{t}^{\theta}(x_{1}\mid x)$ to approximate the unknown posterior $p_{t}(x_{1}\mid x)$, where the learned velocity $v_{t}^{\theta}$ is expressed as the expectation of the conditional velocity under this variational approximation over trajectories. Then, the VFM objective is to minimize the KL divergence between joint distributions, i.e.:
@@
$\mathcal{L}_{\mathrm{VFM}}(\theta)=-\mathbb{E}_{t,x_{1},x}\left[\sum_{d=1}^{D}\log q_{t}^{\theta}(x_{1}^{d}\mid x)\right],\text{ e.g. }\mathcal{L}_{\mathrm{VFM}}(\theta)=\mathbb{E}_{t,x_{1},x}\left[\,\|\mu_{t}^{\theta}(x)-x_{1}\|^{2}\right]$ (5)
in case $q_{t}^{\theta}$ is Gaussian, relating VFM directly back to FM (see Eijkelboom et al. (2024) for details). For sampling with the standard flow matching case of linear interpolation, the vector field reduces to the first moment of the variational approximation:
$v_{t}^{\theta}(x)=\mathbb{E}_{q_{t}^{\theta}(x_{1}|x)}\left[\frac{x_{1}-x}{1-t}\right]=\frac{\mathbb{E}_{q_{t}^{\theta}(x_{1}|x)}[x_{1}]-x}{1-t}=\frac{\mu_{t}^{\theta}(x)-x}{1-t}.$ (6)
A key feature of VFM is its flexibility in choosing $q_{t}^{\theta}$, as different choices allow adaptation to various geometries and data types, improving efficiency and expressiveness.
@@ 3 Riemannian Gaussian Variational Flow Matching
  • Flexibility on the support of the distribution: the prior $p_{0}$ can be defined either on $\mathcal{M}$ (intrinsic) or in the ambient Euclidean space (extrinsic), while vanilla RFM only supports the intrinsic viewpoint. The extrinsic framework maintains the simplicity and efficiency of a linear flow in Euclidean space, avoiding the need for the manifold’s exponential and logarithmic maps, while encoding more geometric information than purely Euclidean methods (section 3.1). Note that, while intrinsic and extrinsic traditionally refer to the manifold’s internal geometry versus its embedding, we use these terms to distinguish whether points lie on the manifold or in the ambient space, rather than coordinate choices. For example, our intrinsic framework can be expressed using ambient coordinates.
@@ 3.1 The Riemannian Gaussian VFM Objective
To extend VFM to the geometric case, one first needs to define a relevant variational posterior with support over the manifold. In contrast to Euclidean settings, we need to take particular care to ensure the distribution is properly defined on the manifold. Let $\mathcal{M}$ be a Riemannian manifold with metric $\mathbf{g}$: the Riemannian Gaussian (RG) distribution (Pennec, 2006) is defined as the maximum–entropy distribution specified by its mean value and covariance, formally
@@
In the Euclidean Gaussian VFM case, this setting reduces to a straightforward mean squared error optimization, so it is natural to wonder whether a similar simplification holds here. In fact, such a simplification exists under two assumptions: (1) the manifold is homogeneous – that is, any point can be transformed into any other by a distance-preserving symmetry (a formal definition is provided in section C.1); and (2) we have access to a closed-form expression for its geodesics. Notably, these requirements are not too restrictive, as most manifolds used in deep learning satisfy them, including $S^{n}$, $\mathbb{H}^{n}$, $\mathbb{T}^{n}$, and $SO(n)$. Formally, the following holds (see section D.1 for details):
@@
Algorithm 1 RG-VFM intrinsic
0:  base $p\in{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\mathcal{M}}$, target $q\in\mathcal{M}$.
 Initialize parameters $\theta$ of $\mu_{t}$
#\# Training Phase
while not converged do
  sample $t\sim{\mathcal{U}}(0,1)$, $x_{0}\sim p,x_{1}\sim q$
[…]
  $\ell(\theta)=\mathbb{E}_{t,x_{1},x}\left[{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\text{dist}_{g}^{2}(x_{1},\mu_{t}(x_{t};\theta))}\right]$
  $\theta=\texttt{optimizer\_step}(\ell(\theta))$
end while
#\# Generation Phase
 sample noise $x_{0}\sim p$
[…]
Algorithm 2 RG-VFM extrinsic
0:  base $p\in{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbb{R}^{d}}$, target $q\in\mathcal{M}$.
 Initialize parameters $\theta$ of $\mu_{t}$
#\# Training Phase
while not converged do
  sample $t\sim{\mathcal{U}}(0,1),x_{0}\sim p,x_{1}\sim q$
[…]
  $\ell(\theta)=\mathbb{E}_{t,x_{1},x}\left[{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\text{dist}_{g}^{2}(x_{1},\mu_{t}(x_{t};\theta))}\right]$
  $\theta=\texttt{optimizer\_step}(\ell(\theta))$
end while
#\# Generation Phase
 sample noise $x_{0}\sim p$
[…]
Minimizing this loss is equivalent to computing the Fréchet mean of the distribution, that is: $\mu^{\star}=\operatorname*{arg\,min}_{\mu_{\theta}\in\mathcal{M}}\mathbb{E}_{x_{1}}\left[\text{dist}_{\mathbf{g}}(x_{1},\mu_{\theta})^{2}\right]$, averaged over the generative steps $t$ and samples $x$. In other words, the point $\mu_{\theta}$ minimizes the expected squared geodesic distance to the target (Fréchet, 1948). Intuitively, this can be viewed as a generalization of the mean squared error from the Euclidean setting to a Riemannian framework. We obtain this result by assuming that $\sigma_{t}(x)$ is constant. Nonetheless, this term could for example be set to $\sigma_{t}(x)=1-t$ to achieve time normalization, as done in our material and protein generation experiments (sections 5.2 and 5.3).
@@
The extrinsic variant RG-VFM-$\mathbb{R}^{n}$ thus learns a simple linear flow while retaining a geometry-aware loss, whereas the intrinsic variant RG-VFM-$\mathcal{M}$ mirrors the RFM setup but differs in its loss definition (algorithms 1 and 2). Indeed, Vanilla RFM also requires $\operatorname{supp}(p_{0})\subseteq\mathcal{M}$ because its vector fields depend on the manifold’s intrinsic geometry. Because of these different frameworks, direct comparison is only meaningful between RG-VFM-$\mathcal{M}$ and RFM, a comparison we present in the next section. The choice between intrinsic and extrinsic versions represents a trade-off: the extrinsic version can only be used in an ambient space $\mathbb{R}^{d}$ of sufficiently large dimension to embed the manifold without degeneracy. In such cases, linear interpolation simplifies implementation and reduces computational costs by requiring only the geodesic distance, rather than logarithm and exponential maps at every step.
@@ 4.1 Jacobi Field formulation of the Flow Matching Objectives
Borrowing the notations from (Chen and Lipman, 2024), we denote the target velocity $v^{0}=u_{t}(x\mid x_{1})$, the predicted velocity $v^{1}=v_{t}^{\theta}(x)$, and their respective endpoints $\gamma_{0}(1)=x_{1}$ and $\gamma_{1}(1)=\mu_{t}^{\theta}(x)$. The losses can be formulated in the Jacobi field framework with the following:
@@
Consider a Jacobi field $J(\tau):=\partial_{s}\alpha(s,\tau)\big|_{s=0}$ as defined in definition 4.1. We denote ${\mathcal{L}}_{\mathrm{RFM}}$ the loss function of the (vanilla) Riemannian Flow Matching (Chen and Lipman, 2024), and ${\mathcal{L}}_{\mathrm{RG\text{-}VFM}}$ the loss function for our proposed Riemannian Variational Flow Matching. Then the following equalities hold:
@@ 4.2 Relation between RG-VFM and RFM Objectives
The curvature functional $\mathcal{C}$ captures how the manifold’s geometry affects the loss comparison, encoding the first- and second-order effects of curvature on geodesic deviation. Thus, RG-VFM implicitly captures the full geometric structure through the exact Jacobi field $J(1)$, while RFM uses only the linear approximation $D_{\tau}J(0)$. This lack of curvature information results in weaker, less precise supervision in directing the flow toward the actual endpoint, leading in practice to RG-VFM learning the signal more effectively than RFM. Special cases are:
@@
  • In spaces of constant curvature (e.g. hyperspheres or hyperbolic spaces) $\nabla R=0$. In this setting, we can restate the result of proposition 4.3 in terms of the constant sectional curvature $K$. The formulation and proof are given in corollary D.1, and in the experimental section we focus primarily on manifolds that fall within this category.
In summary, we introduced RG-VFM as an alternative to RFM for learning a velocity field on a manifold, providing a variational formulation whose objective fully captures higher-order curvature effects, unlike RFM. This results in generally different objectives on curved manifolds. In Euclidean space, however, the RFM objective reduces to CFM, while RG-VFM reduces to VFM (assuming a Euclidean Gaussian posterior rather than Riemannian), and the CFM and VFM objectives become equivalent under appropriate normalization. These relations are schematized in fig. 1 and fig. 2(b), and their schematic algorithms can be compared in section D.2. In terms of computational costs, extrinsic RG-VFM has the same complexity as VFM during both training and sampling. The only difference between the two methods is that VFM compares endpoints using Euclidean distance, while we use geodesic distance. Since we assume geodesic distance to be in closed-form, this introduces no additional computational overhead compared to VFM. Similarly, implicit RG-VFM maintains the same complexity as RFM, with the main difference being that velocity computation happens during sampling rather than during training (see algorithms 3, 4, 5, 6, 8 and 7).
@@ 5.1 Curvature Effects in Synthetic Data
We conduct two sets of experiments: we (1) compare the extrinsic models in their ability to capture the correct geometry – assessed by the distance of the generated samples to the data manifold – and (2) evaluate vanilla versus variational models in reproducing the target distribution. For this last point, the evaluation is based on metrics such as Coverage (% of generated points falling within the desired checkerboard area) and Classifier 2-Sample Tests (C2ST) metric (Lopez-Paz and Oquab, 2016; Dalmasso et al., 2020; Lueckmann et al., 2021). The C2ST technique utilizes a neural network classifier to separate true samples from generated ones, where a score of 0.5 indicates the distributions are indistinguishable to the classifier, while scores approaching 1 suggest the distributions are easily separable. Additional experimental details are provided in appendix E.